Simplify: 30/12
a. 30/12 b. 5/2 c. 10/3 d. 10/4 e. 7/3
Simplify: 13/39
a. 3 b. 13/39 c. 1/9 d. 1/26 e. 1/3
Multiply: 8/9 * 27/2
a. 243/16 b. 16/243 c. 1/12 d. 16/3 e. 12

Answer by jessica43(140)

a) Simplifying (or reducing) fractions means to make the fraction as simple as possible. To do this, try dividing both the top and bottom of the fraction until you can't go any further. Or you can divide both the top and bottom of the fraction by the Greatest Common Factor.
In this case, both 30 and 12 can be divided by 3.
30/12=
(30/3)/(12/3)=
10/4
Now you can divide the both 10 and 4 by 2
10/4=
(10/2)/(4/2)=
5/2
The answer is b) 5/2.
b)13/39
Since 13 is a prime number, that means it can only be divided by 1 and 13. So try dividing 39 by 13 to see if this fraction can be reduced.
13/39=
(13/13)=(39/13)=
1/3
The answer to is e) 1/3.
c)(8/9)*(27/2)
To multiply fractions, you follow 3 steps:
1. Multiply the top numbers (the numerators)
2. Multiply the bottom numbers (the denominators)
3. Simplify
(8/9)*(27/2)=
(8*27)/(9*2)=
216/18
Now you need reduce the fraction as we have done in parts a and b.
216/18=
(216/18)/(18/18)=
12/1=
12
The answer is e) 12.